RON started a business on 1 July2019. He brought in the following assets:
Motor vehicle RM68,000
Cash RM50,000
Equipment and fittings RM35,000
Inventory RM42,000
On 9 August2019,Ronintroducedadditional cash of RM28,000.
After six months oftrading,a net profit of RM25,000was earned.
During thesix months, Ronwithdrew RM18,000 for his personal use.
Required:
(i)What is Ron’sopening capital as at 1 July2019?
(ii)What is Ron’s capital as at 31 December2019?
1
Expert's answer
2020-04-26T19:02:35-0400

(i)Ron’s opening capital as at 1 July 2019 is:

68 000+50 00+35 000+42 00=195 000

this is opening capital


(ii) Ron’s capital as at 31 December 2019 is:

opening capital: 195 000+28 000=223 000

net profit: 25 000

Total capital:223 000+25 00=248 000